OpenStudy (anonymous):

Simplify. x^2 - 36 over 42 x - 7 x^2

OpenStudy (anonymous):

Is this your question:\[\frac{x^2-36}{42x-7x^2}\]

OpenStudy (anonymous):

yes

OpenStudy (lgbasallote):

\[\huge \frac{x^2 - 36}{42x - 7x^2} \implies \frac{(x+6)(x-6)}{7x(6 - x)}\]

OpenStudy (lgbasallote):

i just factored it out

OpenStudy (lgbasallote):

do you know how to solve the problem now??

OpenStudy (anonymous):

no

OpenStudy (anonymous):

your like a lightning bolt @lgbasallote

OpenStudy (lgbasallote):

remember that (a-b) = -(b-a) therefore \[\huge \frac{(x+6)(x-6)}{7x(6 - x)} \implies \frac{(x+6)(x-6)}{-7x(x-6)}\] now do you know how to solve it?

OpenStudy (anonymous):

I have no idea

OpenStudy (anonymous):

hint find the number in the top and bottom then cancel out or get rid off after that u know how to solve it

OpenStudy (lgbasallote):

remember that \[\huge \frac{ab}{ad} \implies \frac{\cancel a b}{\cancel a d} \implies \frac bd\]

OpenStudy (anonymous):

that look the same